OpenStudy (anonymous):

if f(x)=2x-3 & g(x)=x, What is f(g(2))

OpenStudy (hba):

what will be g(2)=?

OpenStudy (hba):

Just replace x with 2.

OpenStudy (anonymous):

-2?

OpenStudy (hba):

How come ? g(x)=x

OpenStudy (hba):

so g(2)=

OpenStudy (anonymous):

oh is it 0?

OpenStudy (hba):

if, g(x)=x then g(2)=2

OpenStudy (ghazi):

\[f(g(x))=2g(x)-3=2*x-3=2x-3\]

OpenStudy (hba):

Now just find f(2)or f(g(2)=?

OpenStudy (ghazi):

\[f(g(2))=2*2-3=4-3\] put x=2

OpenStudy (anonymous):

x cant equal 2 .

OpenStudy (hba):

How can that be ?

OpenStudy (ghazi):

we are supposed to find f(g(2)) which means value of f(g(x)) when x=2

OpenStudy (anonymous):

my choices are -2,0,1,4

OpenStudy (ghazi):

4-3=?

OpenStudy (ghazi):

@Katieann16 did you get it now?

OpenStudy (anonymous):

1

OpenStudy (hba):

@Katieann16 First try to understand how to do it,the answer is 1.But do you understand how to do it or were you just into the options ?

OpenStudy (ghazi):

correct

OpenStudy (anonymous):

i understand now

OpenStudy (ghazi):

good :)

OpenStudy (hba):

You were just so much into the options,what can i say that you did not realize when i asked you g(2)=?

OpenStudy (hba):

@Katieann16 Please tell me if you do not understand ?